Question 26 1 pts

Our company signed a 60-day 6% note for $20,000. Using a 360-day year, what is the total interest due on the maturity date?

$100
$200
$400
$1,200

Flag this Question

Question 27 1 pts

Our company has an account receivable for $12,500 that we have now deemed uncollectible. We use the direct write-off method. Which of the following accounts would we credit to record the write-off?

accounts receivable
allowance for doubtful accounts
bad debt expense
cash

Flag this Question

Question 28 1 pts

Our company uses the percentage of sales method to estimate bad debt expense for the year. Our allowance for bad debts account has a credit balance of $1,000 prior to the adjusting entry for bad debt expense. We have estimated that 2% of net credit sales will be uncollectible for the current year. Net credit sales for the year totaled $200,000. What will be the balance in allowance for bad debts after the adjusting entry is recorded?

$3,000
$4,000
$5,000
$6,000

Flag this Question

Question 29 1 pts

Our company uses the percentage of receivables method to estimate bad debt expense for the year. We had the following account balances on our unadjusted trial balance at the end of the year (December 31): accounts receivable, debit balance of $150,000; allowance for bad debts, credit balance of $1,000. We estimate that 3.5% of accounts receivable at the end of the year are uncollectible. What amount will be debited to bad debt expense when we record the adjusting entry?

$4,000
$4,250
$5,250
$6,250

Flag this Question

Question 30 1 pts

Which method to estimate bad debts is required by GAAP for most companies?

allowance method
direct write-off method
units of production method
FIFO method

Flag this Question

Question 31 1 pts

On January 1, our company purchased a truck for $80,000. The estimated useful life of the truck is 4 years. The residual value at the end of 4 years is estimated to be $10,000. What is the depreciation expense for the third year of use if we use the straight-line method?

$17,500
$20,000
$35,000
$52,500

Please answer all midterm exam thank you
Answer #1

Question - 26= Total Interest = $20,000 × 6% × 60/360 = $200.

Question - 27- Credit the record of write off to Account Receivabel. (Bad debt expense Dr , Acc receivable Cr).

Question - 28 - Bad debt expenses = $200,000×2% = $4,000+$1000 = $5,000.

Question - 29 - Bad debt expense = $150,000 × 3.5% = $5250 - $1000 = $4,250.

Question - 30 -Direct write off method is used by most of the companies.

Question - 31 - Depreciation = (Cost of asset - Salvage value)/ Useful life.

= ($80,000 - $10,0000)/4 = $17,500 p.a

Depreciation for 3rs year = $17,500.